# Comparing Computational Pathology Foundation Models using Representational Similarity Analysis

### v1

@@ 4.4 Effects of Stain Normalization
Given the observed high slide-specificity, we examined whether stain normalization – absent from the default processing of all models – would influence the results. Each patch was normalized using Macenko normalization (Macenko et al., 2009) using the TiaToolbox (Pocock et al., 2022), and the analysis steps were repeated using the normalized patches. Normalization had little effect on the RDM correlation structure between models (Figure 5 and Table 4 in Appendix). However, the slide-specificity of the representations decreased for each model (Table 5 in Appendix). The largest effect was observed for PLIP (20.5%), with CONCH showing the smallest effect (5.5% decrease). Virchow2 showed the lowest slide-specificity after normalization, with Prov-Gigapath remaining the highest. The disease-specificity of the representations also decreased with normalization, though to a lesser degree in magnitude.
@@ 5 Discussion
Beyond assessing performance in downstream tasks, it is important to investigate the representations learned by foundation models. Using techniques popularized in computational neuroscience, we compared the representations of six CPath models and a baseline natural image model across 50,000 image patches and four cancer types. We find that UNI2 and Virchow2 are the most distinct in their representations, as quantified by the average similarity between model RDMs. Having the same pre-training scheme did not guarantee higher similarity between models, as the model most similar to CONCH and KEEP (both vision-language) was Prov-GigaPath (vision-only) and the model most dissimilar to Virchow2 was UNI2 (both vision-only). The causes underlying these findings are unclear, but could potentially relate to differences in datasets and hyperparameters. Future work could systematically analyze how variations in these parameters affect representational structure.
@@
Using the distances between representations, we also quantified the degree of slide- and disease-specificity of each model. All of the CPath models demonstrated a high slide-specificity, where embeddings from patches from the same slide tended to be closer together than embeddings from different slides. This is partly expected and could have benefits for certain tasks – some morphological features would be expected to vary more across tissue samples than within tissue samples, and capturing these variations could help in some downstream precision medicine tasks. However, the reduction in slide-specificity caused by stain normalization also suggests some lack of robustness to slide-specific confounders. Conversely, the models demonstrated relatively low disease specificity. As foundation models have generally shown high downstream performance in tumor type classification (Wang et al., 2024; Bareja et al., 2025), this result is potentially surprising. However, these findings are not necessarily at odds, where certain combinations of features may remain stable across patches/WSIs for a given tumor type, even if there are variations in the representations as a whole.
@@
In terms of intrinsic dimensionality, we found that the vision-language models had relatively lower dimensional representations, whereas the vision-only models had relatively distributed representations. One potential explanation relates to differences in training objectives between these models. Vision-language models are trained to learn image embeddings that are aligned with language embeddings, where this language bottleneck may may inherently encourage compression into lower-dimensional spaces. Conversely, current vision-only models use image-based objectives based on masking and data augmentation, which may promote the preservation of richer visual details. This larger representational capacity may help explain the generally higher performance of vision-only models across collections of downstream tasks (Marza et al., 2025; Bareja et al., 2025), though the higher amounts of training data typically used for vision-only models may also explain this performance.
@@
Our findings have several implications for the development and deployment of foundation models in computational pathology. The observed high slide-specificity and effects of stain normalization suggests some lack of robustness in current models, consistent with other recent studies (Kömen et al., 2024; de Jong et al., 2025). Using data augmentation and/or adversarial learning during model development could potentially improve this robustness (Ganin et al., 2016; Su et al., 2025), and stain normalization may help at inference time. The existence of slide specific representational signatures is not necessarily detrimental to downstream tasks, but it does raise the possibility of shortcut learning and reduced generalization when fitting classifiers on top of the embeddings. Additionally, our findings regarding the pairwise similarities between models can inform ensembling strategies. Ensembling of different models often improves downstream performance, yet the high computational costs of CPath foundation models can limit the number of models that can be ensembled. Thus, ensembling could potentially focus on more dissimilar, complementary models.
@@
Our study has several limitations. While we studied four disease cohorts, all of these cohorts came from the TCGA dataset, so the results may quantitatively differ in different datasets. More broadly, it will be valuable in future work to explicitly assess generalizability across different institutions, staining protocols, and scanner hardware, including how the studied representational metrics relate to downstream performance. Another challenge of RSA is its high CPU memory demand, which scales quadratically with the number of samples ($N^{2}$) when computing representational dissimilarity matrices. This places constraints on sample size and can require high-memory computing nodes.

### v2

@@ 4.4 Effects of Stain Normalization
Given the observed high slide-specificity, we examined whether stain normalization – absent from the default processing of all models – would influence the results. Each patch was normalized using Macenko normalization (macenko) using the TiaToolbox (tiatoolbox), and the analysis steps were repeated using the normalized patches. Normalization had little effect on the RDM correlation structure between models (Figure 5 and Table 4 in Appendix). However, the slide-specificity of the representations decreased for each model (Table 7 in Appendix). The largest effect was observed for PLIP (20.5%), with CONCH showing the smallest effect (5.5% decrease). Virchow2 showed the lowest slide-specificity after normalization, with Prov-Gigapath remaining the highest. The disease-specificity of the representations also decreased with normalization, though to a lesser degree in magnitude.
4.5 Hyperparameter Sensitivity
To assess robustness to the patch sampling configuration, we repeated all analyses using a different number of patches per slide. The original results were computed over five independent batches of 50 WSIs × 50 patches/WSI (× 4 diseases = 10,000 images). The new sampling configuration used five independent batches of 25 WSIs × 100 patches/WSI (× 4 diseases = 10,000 images). We observed highly similar results: the Pearson correlation was 0.997 between the original model-similarity values (Figure 2) and the new values (Figure 6), with UNI2 and Virchow2 again exhibiting the lowest average correlations and Prov-GigaPath the highest. The Pearson correlations for slide-specificity and disease-specificity were 0.999 and 0.984, respectively (Table 3, Table 8).
The original analyses used Euclidean distance for RDM construction, recommended for its robustness to representational noise (Botero2024; Bosch2025). Because several models were trained with contrastive loss, we additionally evaluated whether using a correlation-based metric would alter the RSA results. Using Pearson correlation for RDM construction yielded similar findings: Virchow2 again exhibited the lowest average similarity across models and Prov-GigaPath the highest. The Pearson correlation between the full set of original similarity values between models (Figure 2) and the new values (Figure 7) was 0.84, indicating high correlation with some expected quantitative variation. In particular, UNI2’s average similarity between models increased from 0.421 to 0.527. Thus, while there are some quantitative changes with a different distance metric, the overall findings remain robust. We note that the additional experiments (slide/disease specificity, spectral analysis) are independent of the RDM construction.
@@ 5 Discussion
Beyond assessing performance in downstream tasks, it is important to investigate the representations learned by foundation models, which can offer insights into the strengths and weaknesses of current approaches and the influence of training strategies on embedding structure. Using techniques popularized in computational neuroscience, we compared the representations of six CPath models and a baseline natural image model across 50,000 image patches and four cancer types.
@@
We find that UNI2 and Virchow2 are the most distinct in their representations, as quantified by the average similarity between model RDMs. Having the same pre-training scheme did not guarantee higher similarity between models, as the model most similar to CONCH and KEEP (both vision-language) was Prov-GigaPath (vision-only) and the model most dissimilar to Virchow2 was UNI2 (both vision-only). The causes underlying these findings are unclear, but could potentially relate to differences in datasets or hyperparameters. Future work could systematically analyze how variations in these parameters affect representational structure.
One notable observation is the higher correlation between PLIP and the natural image baseline (ViT-DINOv2). One possible explanation is that PLIP was trained on fewer pathology images (208K) than the other models (each over 1M). As a result, its representations may partially overlap with the broader visual embeddings of natural images. Nevertheless, this does not preclude strong performance in certain tasks, such as PLIP performing well on zero-shot classification in external datasets (huang2023visual), highlighting that representational overlap with natural images can still support robust transfer in clinically relevant tasks.
@@
Using the distances between representations, we also quantified the degree of slide- and disease-specificity of each model. All of the CPath models demonstrated a high slide-specificity, where embeddings from patches from the same slide tended to be closer together than embeddings from different slides. This is partly expected and could be beneficial for certain tasks – some morphological features would be expected to vary more across tissue samples than within tissue samples, and capturing these variations could be important in some precision medicine tasks. However, the reduction in slide-specificity caused by stain normalization also suggests some lack of robustness to slide-specific confounders. Conversely, the models demonstrated relatively low disease specificity. As foundation models have generally shown high downstream performance in tumor type classification (Wang2024-qz; Bareja2025), this result is potentially surprising. However, these findings are not necessarily at odds, where certain combinations of features may remain stable across patches/WSIs for a given tumor type, even if there are variations in the representations as a whole.
@@
In terms of intrinsic dimensionality, we found that the vision-language models had relatively lower-dimensional representations compared to the relatively distributed representations of vision-only models. One potential explanation relates to differences in training objectives between these approaches. Vision-language models are trained to align image embeddings with corresponding language embeddings, introducing a language bottleneck wherein visual representations are forced to align with lower-dimensional, semantically structured language embeddings in a shared multimodal space. This alignment may thus encourage compression of rich visual details into more compact representations. In contrast, current vision-only models use image-based objectives based on masking and data augmentation, which may promote the preservation of richer visual details.
Another interesting observation was that Prov-GigaPath exhibited high representational similarity across models, high slide specificity, and high intrinsic dimensionality These observations are not mutually exclusive and may, in fact, be related. For example, because all models demonstrate high slide specificity, a model with the greatest slide specificity may then show higher representational similarity to other models. Additionally, higher intrinsic dimensionality may correlate with capturing more fine-grained, slide-specific features. Consistent with this interpretation, the lowest intrinsic dimensionality was observed for CONCH, which also exhibited the lowest slide specificity, followed by Virchow2, which had the second lowest values for both.
@@
Our findings have several implications for the development and deployment of foundation models in computational pathology. The observed high slide-specificity and effects of stain normalization suggests some lack of robustness in current models, consistent with other recent studies (komen2024; dejong2025). Using data augmentation and/or adversarial learning during model development could potentially improve this robustness (dann; advdino), and stain normalization may help at inference time. The existence of slide-specific representational signatures is not necessarily detrimental to downstream tasks, but it does raise the possibility of shortcut learning and reduced generalization when fitting classifiers on top of the embeddings. Additionally, our findings regarding the pairwise similarities between models can inform ensembling strategies. Ensembling of different models often improves downstream performance, yet the high computational costs of CPath foundation models limits the number of models that can be ensembled. Thus, ensembling could focus on more dissimilar, complementary models.
Several recent works have provided extensive downstream evaluation of pathology foundation models (marza2025; Bareja2025). One intriguing potential connection between those results and our findings relates to Virchow2. Bareja2025 found that vision-only models outperformed other model types and Virchow2 had the highest overall performance of all models. In our study, we also found that Virchow2 had lower slide specificity than the other vision-only models. While this is not a causal connection, it suggests potential links to motivate follow-up work, where our study provides a unique and complementary perspective compared to other efforts in the field.
@@
While we studied four disease cohorts, all of these cohorts came from TCGA, so the results may quantitatively differ in different datasets. More broadly, it will be valuable in future work to explicitly assess generalizability across different institutions, staining protocols, and scanner hardware, including how the studied representational metrics relate to downstream performance. Another challenge of RSA is its high CPU memory demand, which scales quadratically with the number of samples ($N^{2}$). This places constraints on sample size and can require high-memory computing nodes.
@@ Conclusion
[…]
W.L. acknowledges funding support from the Ellison Foundation, the Wong Family Award, the Louis B. Mayer Foundation, the National Institute of Biomedical Imaging and Bioengineering award R21EB035247, and the National Library of Medicine award R01LM014775.
@@
Figure 6: Spearman correlation between the RDMs of each pair of models using 25 WSIs per disease with 100 patches per WSI. Analogous to Figure 2 except using the modified sampling scheme.
Figure 7: Spearman correlation between the RDMs of each pair of models using Pearson correlation distance for RDM construction instead of Euclidean. Analogous to Figure 2 except using the different distance metric.
@@
Table 5: Mean Spearman correlation between the RDM for each model and the other five CPath models using 25 WSIs per disease with 100 patches per WSI. Analogous to Table 2 except using this modified sampling scheme.
Model Pretraining Mean Corr.
[…]
Virchow2 0.423
Prov-Gigapath 0.525
CONCH Vision-Language 0.463
PLIP 0.442
KEEP 0.452
Table 6: Mean Spearman correlation distance between the RDM for each model and the other five CPath models when using Pearson correlation distance for RDM construction. Analogous to Table 2 except using the different distance metric.
Model Pretraining Mean Corr.
UNI2 Vision 0.527
[…]
Prov-Gigapath 0.578
CONCH Vision-Language 0.472
PLIP 0.481
KEEP 0.472
@@
Table 8: Assessment of slide- and disease-specificity of model representations when sampling 25 WSIs per disease with 100 patches per WSI. Analogous to Table 3 except using the different sampling parameters.
Slide Specificity Disease Specificity
Model Cliff’s Delta Range Cliff’s Delta Range
UNI2 0.773 [0.742, 0.804] 0.196 [0.163, 0.216]
[…]
Prov-Gigapath 0.795 [0.778, 0.808] 0.184 [0.159, 0.205]
[…]
PLIP 0.730 [0.710, 0.745] 0.190 [0.129, 0.236]
KEEP 0.728 [0.698, 0.750] 0.245 [0.240, 0.252]
ViT-Dinov2 0.388 [0.336, 0.415] 0.088 [0.070, 0.099]